OpenStudy (anonymous):

i have a math problem that is for estimating 6 3/4 + 7 3/4 which i get it estimated to 7 + 8 which gives me 15 but the back of the book says its 14 1/2 can you help me with this problem

jhonyy9 (jhonyy9):

6 3/4 = 27/4 7 3/4 = 31/4 27/4 + 31/4 = 58/4 = 14 2/4 = 14 1/2